Problem 3 A single conduction loop of wire has an area of 7.4x102 m2 and a resistance of 1102. Perpendicular to the plane of the loop is a magnetic field of strength 0.18T. At what rate (in T/s) must this field change if the induced current in the loop is to be 0.22A? search
Answer #1

Voltage = I*R

= 0.22*110

= 24.2 V

V = A dB/dt

24.2 = 0.074*dB/dt

dB/dt = 327.03 T/s

so the rate of change of magnetic field = 327.03 T/s